# Generalized Momenta-Based Koopman Formalism for Robust Control of Euler-Lagrangian Systems

### v1

@@ I Introduction
  • We reformulate the Euler-Lagrangian dynamics in an implicit momentum-based representation that decouples the control channel from state-dependent dynamics, thereby yielding a structure naturally suited for learning linear Koopman models under actuation.
  • The linear model learned from the proposed approach exhibits a superior prediction performance compared to the more computationally demanding bilinear models learned using the conventional explicit representation of the dynamics.
@@
  • We validate the proposed approach through extensive simulation and experimental studies to demonstrate the superior tracking accuracy and disturbance rejection compared to state-of-the-art methods.
The remainder of the paper is organized as follows. Section II presents the preliminaries for Koopman operator theory. Sections III and IV present the proposed momentum-based Koopman formulation and neural architectures employed for learning the subsequent Koopman models. Sections V and VI detail GESO for disturbance estimation with its subsequent integration within a MPC framework. Section VII describes simulation and experimental results. Section VIII concludes the paper.
@@ II Preliminaries
where $u_{i}\in\mathbb{R}^{m}\;\forall\;i=1,2,...,m$ denotes the control inputs. $\bm{f}_{0}:\mathbb{X}\to\mathbb{X}$ and $\bm{f}_{i}:\mathbb{R}^{m}\to\mathbb{X}\;\forall\;i=1,2,...,m$ represent the drift and state-dependent control vector fields, respectively. The corresponding Koopman representation takes the form:
@@
III Generalized Momentum-based Koopman Formulation
@@
where $\bm{q}\in\mathbb{R}^{n}$ and $\bm{\tau}\in\mathbb{R}^{m}$ refer to the position coordinates and control inputs, respectively. $\bm{M}(\bm{q})\in\mathbb{R}^{n\times n}$, $\bm{C}(\bm{q},\,\bm{\dot{q}})\in\mathbb{R}^{n}$ and $\bm{G}(\bm{q})\in\mathbb{R}^{n}$ represent the mass matrix, the Coriolis terms and the gravity term, respectively. Traditionally, position ($\bm{q}$) and velocity ($\bm{\dot{q}}$) coordinates are chosen as generalized coordinates to explicitly represent the state space for the Euler-Lagrangian system. The evolution of the generalized velocities is given as
@@
Due to their interaction with the mass matrix, control vectors become state-dependent. Consequently, this explicit representation lends itself more naturally to bilinear formulation (4) [10]. Although expressive, bilinear Koopman models are computationally more expensive, require larger datasets, and often preclude integration with non-linear MPC frameworks [12].
@@
$\bm{x}=\begin{bmatrix}\bm{q}\\ \bm{M}(\bm{q})\bm{\dot{q}}\end{bmatrix}=\begin{bmatrix}\bm{x}_{u}\\ \bm{x}_{a}\end{bmatrix}$, (8)
where $\bm{x}_{u}$ and $\bm{x}_{a}$ represent the actuated and the unactuated part of the dynamics, respectively. The evolution of the proposed state space vector is given as
@@
where $\bm{f}(\bm{x})=[\bm{f}_{u}(\bm{x}),\bm{f}_{a}(\bm{x})]^{\top}$ and $\bm{\tilde{u}}=[\bm{0}_{n},\bm{u}]^{\top}$ where $\bm{0}_{n}\in\mathbb{R}^{n}$ represents the zero vector and $\bm{u}=\bm{\tau}$. Hence, the proposed choice of state space completely decouples the linear control inputs from the state-dependent dynamics, making resultant dynamics more amenable for the formulation of linear Koopman models (II).
@@
Consider the Koopman lifted state to be $\bm{z}=[\bm{x}^{\top},\bm{\phi}(\bm{x})^{\top}]^{\top}$ where $\bm{\phi}(\cdot):\mathbb{R}^{n}\to\mathbb{R}^{N}$ represents the observable functions. Then, as per [13], we can construct a Koopman matrix to relate the passive and the actuated dynamics as
@@
where $\bm{A}_{c}=\begin{bmatrix}\bm{A}_{c,xx}&{A}_{c,x\phi}\\ \bm{A}_{c,x\phi}&{A}_{c,\phi\phi}\end{bmatrix}$, $\bm{B}_{c}=\begin{bmatrix}\bm{0}_{n\times m}\\ \bm{I}_{n\times m}\\ \bm{0}_{N\times m}\end{bmatrix}$, $\bm{C}^{x}=\begin{bmatrix}\bm{I}_{n\times n}\\ \bm{0}_{n\times N}\end{bmatrix}$. $\bm{I}_{n\times n}$ represents a $n\times n$ identity matrix. The corresponding discrete time Koopman model is
@@
Figure 1: Framework of the proposed momentum-based Koopman-MPC-GESO architecture. a) Model Training. b) Control framework. c) Trajectory tracking results.
@@ IV Neural Architectures for learning Passive Dynamics
Both networks share a common loss structure designed to ensure accurate lifting and prediction, within the Koopman space. The constituent components of the loss are defined as follows:
@@
Figure 2: Prediction performance comparison for the proposed scheme with Nominal Linear Koopman (NLK) and Nominal Bilinear Koopman (NBK) [12]. Testing errors across a) different manipulators, and b) different sizes of training datasets for the 4R manipulator.
@@ VI Model Predictive Control
The proposed Koopman architecture is paired with linear MPC. Linear MPC with a prediction (and control) horizon $s$ is defined as an optimization problem of the form
@@
where $\bm{x}_{ref,k}$ refers to the reference state at $k^{th}$ timestep. $\bm{R}$ represents the penalty matrix, while $\bm{x}^{-},\,\bm{x}^{+}]$ and $\bm{u}^{-},\,\bm{u}^{+}]$ represent the state and input constraints, respectively. The full proposed momentum-based Koopman-MPC-GESO architecture is shown in Fig. 1.
Figure 3: Tracking performance comparisons for a 3R serial manipulator with a) No disturbance. b) Actuator fault ($\bm{\tau_{d}}=0.6\bm{u}\;\text{Nm}$). c) Constant disturbance torque ($\bm{\tau_{d}}=[10,10,10]\;\text{Nm})$ d) Constant end effector load($\bm{F_{d}}=[20,20,20]\;\text{N}$}).
@@ VII Results
This section presents a comprehensive evaluation of the proposed architecture through simulation and experimental studies on serial manipulators. The assessment considers both open-loop prediction accuracy and closed-loop control performance, thereby providing a rigorous demonstration of the effectiveness and robustness of the proposed framework.
@@ VII-A Prediction Performance Comparison
Testing errors are used as comparison metrics. The testing errors are standardized to ensure fair comparison between the proposed scheme ($\bm{x}=[\bm{q},\bm{M\dot{q}}]^{\top}$) and the baseline nominal scheme ($\bm{x}=[\bm{q},\bm{\dot{q}}]^{\top}$), which employ different state space coordinates. The corresponding performance comparison across different ranges of manipulators and sizes of datasets is shown in Fig. 2. The proposed scheme outperforms both NLK and, more importantly, NBK [12] across all the evaluations. This result is particularly noteworthy because bilinear formulations are typically regarded as better suited for Euler-Lagrangian systems due to the intrinsic coupling between states and inputs. Nevertheless, the momentum-based formulation achieves superior performance by structurally separating the linear input channel from the passive nonlinear dynamics, thereby making the system more amenable to linear Koopman modeling. This structural advantage not only reduces modeling complexity but also enables higher prediction accuracy than bilinear approaches, reconciling the dual objectives of modeling accuracy and computational efficiency in a way that bilinear approaches cannot. Further, it can be observed (Fig. 2b) that the linear models only require 100 trajectories to achieve the optimal performance, with the proposed scheme consistently outperforming the baselines across all dataset sizes. While NBK [12] requires a much larger dataset with about 250 trajectories to achieve convergence in terms of prediction capabilities while still offering considerably inferior performance.
@@ VII-B Closed Loop Control Performance Comparison
Figure 4: Experimental results for trajectory tracking with a) Franka Research 3 manipulator b) without and c) with external disturbances.
@@
Subsequently, we consider the tracking evaluation for trajectory tracking under the influence of various types of disturbance. In particular, we consider three types of disturbances namely, actuator faults ($\bm{d}_{1}$) with $\bm{\tau}_{d}=-0.6\bm{u}\penalty 10000\ Nm$, constant torque disturbance ($\bm{d}_{2}$) with $\bm{\tau}_{d}=[10,10,10]\penalty 10000\ Nm$}, and constant end effector load ($\bm{d}_{3}$) of $\bm{F_{d}}=[20,20,20]\penalty 10000\ N$}. The corresponding results are shown in Figs. 3b - 3d and Table I. As can be seen, even under such extreme external disturbances, the proposed framework is able to track the desired trajectory effectively. This robustness is achieved by the incorporated GESO, which compensates for disturbances in real-time. In contrast, the baseline NLK architecture fails to mitigate the effects of these disturbances, resulting in deteriorated tracking performance and significant deviation from the desired path. Further, the proposed scheme also significantly outperforms KGP [19] and RAL [14] schemes, both of which are equipped with capabilities, a Gaussian process, and higher-order GESOs, respectively, to deal with disturbances. The relatively poor performance of RAL can be attributed to the limited prediction accuracy of the offline Koopman model obtained through active learning, which is inferior to the proposed scheme. Furthermore, since RAL [14] employs higher-order GESOs, these observers cannot be directly embedded within an MPC framework. Instead, they must be appended to the MPC output, necessitating external saturation filters for safety, which acts as an additional hindrance to the overall effectiveness of RAL [14]. In addition, when paired with a high-gain base controller, higher-order GESOs in RAL [14] induce control chattering which introduces discontinuous corrective actions that further degrade closed-loop performance.
@@ VII-C Experimental Results
This subsection evaluates the closed-loop prediction performance of the proposed scheme on FRANKA Research 3 (Fig. 4a) manipulator arm. For learning the Koopman model, 500 trajectories are generated randomly within the workspace of the robot, each consisting of 1000 data points sampled at 200 Hz. The parameters for GESO (V) are $k_{1}=100$ and $k_{2}=2000$. The effect of the proposed algorithms is evaluated by implementing the control architecture to trace a hypotrochoid path. The corresponding results are shown in Fig. 4.
As seen from Fig. 4b and c, the robot is able to trace the desired path accurately both in the presence and absence of the external disturbances, while the NLK model is unable to trace the desired paths in either case. The resulting RMSE errors for the proposed scheme are $0.024\;\text{m}$ and $0.035\;\text{m}$ for the case with and without disturbance, respectively. The corresponding errors for NLK are $0.077\;\text{m}$ and $0.091\;\text{m}$. The poor tracking performance of the NLK stems from a combination of the poor prediction model and a lack of effective disturbance compensation. Hence, the experimental results demonstrate accurate real-time trajectory tracking of the proposed scheme with the added benefit of mitigating disturbances.
@@ VIII Conclusion
This paper presents a momentum-based Koopman operator formulation for Euler-Lagrangian dynamics that employs an implicit generalized momentum–based state space. This representation decouples a known linear actuation channel from the state-dependent dynamics, making the dynamics more amenable to linear Koopman modeling in contrast to explicit state space representations, which lend themselves more naturally to bilinear Koopman formulations. By leveraging this structural separation, the proposed approach requires to learns only the unactuated dynamics rather than the full actuation-dependent system, thereby significantly reducing the number of learnable parameters, improving data efficiency, and lowering overall model complexity. To realize this framework, we introduced two neural network architectures capable of constructing Koopman embeddings from either actuated or unactuated data, enabling flexible deployment under diverse data availability scenarios. Linear Generalized Extended State Observer (GESO) is integrated with proposed scheme to enhance robustness to external disturbances and unmodeled dynamics. Unlike higher-order nonlinear observers employed in previous studies, linear GESOs can be seamlessly incorporated into linear control frameworks, preserving their structure while enhancing stability and robustness. Extensive simulation and experimental studies are conducted to show that the momentum-based Koopman formulation achieves superior predictive accuracy compared to explicit state space representation-based linear and bilinear Koopman models, while the integrated GESO ensures reliable performance under a wide range of disturbances.

### v2

# arXiv:2509.17010v2[cs.RO] 27 Feb 2026

@@ Abstract
Index terms – Koopman operator, Euler–Lagrangian systems, unactuated dynamics, linear Koopman models, generalized extended state observer (GESO), learning-based control
@@ I Introduction
  • We reformulate the Euler-Lagrangian dynamics in an implicit momentum-based representation that decouples control from state-dependent dynamics, enabling the learning of linear Koopman models which exhibit superior prediction performance compared to the more computationally demanding bilinear models learned using conventional explicit representation.
@@
The remainder of the paper is organized as follows. Section II presents the preliminaries for Koopman operator theory. Sections III and IV present the proposed momentum-based Koopman formulation and neural architectures employed for learning the subsequent Koopman models. Sections V and VI detail GESO formulation for disturbance estimation with its subsequent integration within a MPC framework. Section VII describes simulation and experimental results. Section VIII concludes the paper.
@@ II Preliminaries
where $u_{i}\in\mathbb{R}^{m}\;\forall\;i=1,2,...,m$ denotes the control inputs. $\bm{f}_{0}:\mathbb{X}\to\mathbb{X}$ and $\bm{f}_{i}:\mathbb{X}\to\mathbb{X}\;\forall\;i=1,2,...,m$ represent the drift and state-dependent control vector fields, respectively. The corresponding Koopman representation takes the form:
@@
III Generalized Momenta-based Koopman Formulation
@@
where $\bm{q}\in\mathbb{R}^{l}$ and $\bm{\tau}\in\mathbb{R}^{m}$ refer to the position coordinates and control inputs, respectively. $\bm{M}(\bm{q})\in\mathbb{R}^{l\times l}$, $\bm{C}(\bm{q},\,\bm{\dot{q}})\in\mathbb{R}^{l\times l}$ and $\bm{G}(\bm{q})\in\mathbb{R}^{l}$ represent the mass matrix, the Coriolis terms and the gravity term, respectively. Traditionally, position ($\bm{q}$) and velocity ($\bm{\dot{q}}$) coordinates are chosen as generalized coordinates to explicitly represent the state space for the Euler-Lagrangian system. The evolution of the generalized velocities is given as
@@
Due to their interaction with the mass matrix, control vectors become state-dependent. Consequently, this explicit representation lends itself more naturally to bilinear formulation (4) [7]. Although expressive, bilinear Koopman models are computationally more expensive, require larger datasets, and often preclude integration with linear MPC frameworks [2].
@@
$\bm{x}=\begin{bmatrix}\bm{q}\\ \bm{M}(\bm{q})\bm{\dot{q}}\end{bmatrix}=\begin{bmatrix}\bm{x}_{u}\\ \bm{x}_{a}\end{bmatrix}\in\mathbb{R}^{n}$ (8)
where $\bm{x}_{u}$ and $\bm{x}_{a}$ represent the unactuated and the actuated part of the dynamics, respectively. The evolution of the proposed state space vector is given as
@@
where $\bm{f}(\bm{x})=[\bm{f}_{u}(\bm{x}),\bm{f}_{a}(\bm{x})]^{\top}$ and $\bm{\tilde{u}}=[\bm{0}_{l},\bm{u}]^{\top}$ where $\bm{0}_{l}\in\mathbb{R}^{l}$ represents the zero vector and $\bm{u}=\bm{\tau}$. Hence, the proposed choice of state space completely decouples the linear control inputs from the state-dependent dynamics, making resultant dynamics more amenable for the formulation of linear Koopman models (II).
@@
Consider the Koopman lifted state to be $\bm{z}=[\bm{x}^{\top},\bm{\phi}(\bm{x})^{\top}]^{\top}\in\mathbb{R}^{N}$, where $\bm{\phi}(\cdot):\mathbb{R}^{n}\to\mathbb{R}^{P}$ represents the observable functions. Then, as per [1], we can construct a Koopman matrix to relate the passive and the actuated dynamics as
@@
where $\bm{A}_{c}=\begin{bmatrix}\bm{A}_{c,xx}&{A}_{c,x\phi}\\ \bm{A}_{c,x\phi}&{A}_{c,\phi\phi}\end{bmatrix}$, $\bm{B}_{c}=\begin{bmatrix}\bm{0}_{l\times m}\\ \bm{I}_{l\times m}\\ \bm{0}_{P\times m}\end{bmatrix}$, $\bm{C}^{x}=\begin{bmatrix}\bm{I}_{n\times n}&\bm{0}_{n\times P}\end{bmatrix}$. $\bm{I}_{n\times n}$ represents a $n\times n$ identity matrix. The corresponding discrete time Koopman model is
@@
Figure 1: Framework of the proposed momentum-based Koopman-MPC-GESO architecture depicting a) model training, b) control framework, and c) trajectory tracking results.
@@ IV Neural Architectures for learning Passive Dynamics
Both networks share a common loss structure designed to ensure accurate lifting and prediction within the Koopman space. The constituent components of the loss are defined as follows:
@@
((a))
((b))
Figure 2: Predictive performance comparison for the proposed scheme with Nominal Linear Koopman (NLK) and Nominal Bilinear Koopman (NBK) [2]. Testing errors across a) different manipulators, and b) different sizes of training datasets for a 4R manipulator.
@@ VI Model Predictive Control
The proposed Koopman architecture is paired with a linear MPC. Linear MPC with a prediction (and control) horizon $S$ is defined as an optimization problem of the form
@@
where $\bm{x}_{ref,k}$ refers to the reference state at $k^{th}$ timestep. $\bm{R}$ represents the penalty matrix, while $\bm{x}^{-},\,\bm{x}^{+}]$ and $\bm{u}^{-},\,\bm{u}^{+}]$ represent the state and input constraints, respectively. $\dagger$ represents the pseudoinverse. The full proposed momentum-based Koopman-MPC-GESO architecture is shown in Fig. 1.
Figure 3: Tracking performance comparisons for a 3R serial manipulator with a) no disturbance, b) actuator fault ($\bm{\tau_{d}}=-0.6\bm{u}\;\text{Nm}$), c) constant disturbance torque ($\bm{\tau_{d}}=[10,10,10]\;\text{Nm})$, and d) constant end-effector load($\bm{F_{d}}=[20,20,20]\;\text{N}$}).
@@ VII Results
This section presents a comprehensive evaluation of the proposed architecture through simulation and experimental studies on serial manipulators. The assessment considers both open-loop prediction accuracy and closed-loop control performance, thereby providing a rigorous demonstration of the effectiveness and robustness of the proposed framework111code : https://github.com/Rajpal9/Robust_Gen_Mom_Koopman.git.
@@ VII-A Prediction Performance Comparison
Testing errors are used as comparison metrics. The testing errors are standardized to ensure fair comparison between the proposed scheme ($\bm{x}=[\bm{q}^{\top},(\bm{M\dot{q}})^{\top}]^{\top}$) and the baseline nominal scheme ($\bm{x}=[\bm{q}^{\top},\bm{\dot{q}}^{\top}]^{\top}$), which employ different state space coordinates. The corresponding performance comparison across different ranges of manipulators and sizes of datasets is shown in Fig. 2. The proposed scheme outperforms both NLK and, more importantly, NBK [2] across all the evaluations. This result is particularly noteworthy because bilinear formulations are typically regarded as better suited for Euler-Lagrangian systems due to the intrinsic coupling between states and inputs. Nevertheless, the momentum-based formulation achieves superior performance by structurally separating the linear input channel from the passive nonlinear dynamics, thereby making the system more amenable to linear Koopman modeling. This structural advantage not only reduces modeling complexity but also enables higher prediction accuracy than bilinear approaches, reconciling the dual objectives of modeling accuracy and computational efficiency in a way that bilinear approaches cannot. Further, it can be observed (Fig. 2b) that the linear models only require 100 trajectories to achieve the optimal performance, with the proposed scheme consistently outperforming the baselines across all dataset sizes. While NBK [2] requires a much larger dataset with about 250 trajectories to achieve convergence in terms of prediction capabilities while still offering considerably inferior performance.
@@ VII-B Closed Loop Control Performance Comparison
Figure 4: Experimental results for trajectory tracking with the FR 3 manipulator, a,c) without, and b,d) with external disturbances.
@@
Subsequently, we consider the tracking evaluation for trajectory tracking under the influence of various types of disturbance. In particular, we consider three types of disturbances namely, actuator faults ($\bm{d}_{1}$) with $\bm{\tau}_{d}=-0.6\bm{u}\penalty 10000\ Nm$, constant torque disturbance ($\bm{d}_{2}$) with $\bm{\tau}_{d}=[10,10,10]\penalty 10000\ Nm$}, and constant end effector load ($\bm{d}_{3}$) of $\bm{F_{d}}=[20,20,20]\penalty 10000\ N$}. The corresponding results are shown in Figs. 3b - 3d and Table I. As can be seen, even under such extreme external disturbances, the proposed framework is able to track the desired trajectory effectively. This robustness is achieved by the incorporated GESO, which compensates for disturbances in real-time. In contrast, the baseline NLK architecture fails to mitigate the effects of these disturbances, resulting in degraded tracking performance and significant deviation from the desired path. Further, the proposed scheme also significantly outperforms KGP [13] and RAL [14] schemes, both of which are equipped with capabilities, Gaussian processes and higher-order GESOs respectively, to deal with disturbances. The relatively poor performance of RAL [14] can be attributed to the limited prediction accuracy of the offline Koopman model obtained through active learning, which is inferior to the proposed scheme. Furthermore, since RAL [14] employs higher-order GESOs, these observers cannot be directly embedded within an MPC framework. Instead, they must be appended to the MPC output, necessitating external saturation filters for safety, which acts as an additional hindrance to the overall effectiveness of RAL [14]. In addition, when paired with a high-gain base controller, higher-order GESOs in RAL [14] induce control chattering which introduces discontinuous corrective actions that further degrade closed-loop performance.
@@ VII-C Experimental Results
This subsection evaluates the closed-loop prediction performance of the proposed scheme on FRANKA Research 3 manipulator arm. For learning the Koopman model, 500 trajectories are generated randomly within the workspace of the robot, each consisting of 1000 data points sampled at 200 Hz. The parameters for GESO (V) are $k_{1}=100$ and $k_{2}=2000$. The effect of the proposed algorithms is evaluated by implementing the control architecture to trace a hypotrochoid and a tricuspid path.
As seen from Fig. 4, the robot is able to trace the desired path accurately both in the presence and absence of the external disturbances, while the NLK model is unable to trace the desired paths in either case. The resulting RMSE errors for the proposed scheme for the hypotrochoid shape are $0.024\;\text{m}$ and $0.035\;\text{m}$ for the case with and without disturbance, respectively. The corresponding errors for NLK are $0.077\;\text{m}$ and $0.091\;\text{m}$. Similarly for the tricuspid shape, the resulting RMSE errors for the proposed scheme are $0.011\;\text{m}$ and $0.019\;\text{m}$ for the case with and without disturbance, respectively. The corresponding errors for NLK are $0.043\;\text{m}$ and $0.048\;\text{m}$. The poor tracking performance of the NLK stems from a combination of the poor prediction model and a lack of effective disturbance compensation. Hence, the experimental results demonstrate accurate real-time trajectory tracking of the proposed scheme with the added benefit of mitigating disturbances.
@@ VIII Conclusion
This paper presents a momentum-based Koopman operator formulation for Euler-Lagrangian dynamics that employs an implicit generalized momentum–based state space. This representation decouples a known linear actuation channel from the state-dependent dynamics, making the dynamics more amenable to linear Koopman modeling in contrast to explicit state space representations, which lend themselves more naturally to bilinear Koopman formulations. By leveraging this structural separation, the proposed approach requires to learns only the unactuated dynamics rather than the full actuation-dependent system, thereby significantly reducing the number of learnable parameters, improving data efficiency, and lowering overall model complexity. To realize this framework, we introduced two neural network architectures capable of constructing Koopman embeddings from either actuated or unactuated data, enabling flexible deployment under diverse data availability scenarios. Linear Generalized Extended State Observer (GESO) is integrated with proposed scheme to enhance robustness to external disturbances and unmodeled dynamics. Unlike higher-order nonlinear observers employed in previous studies, linear GESOs can be seamlessly incorporated into linear control frameworks, preserving their structure while enhancing stability and robustness. Extensive simulation and experimental studies are conducted to show that the proposed formulation achieves superior predictive accuracy compared to explicit state space representation-based linear and bilinear Koopman models, while the integrated GESO ensures reliable performance under a wide range of disturbances.